List the quadruples in the relation { a,b,c,d} where a, b, c, d are integers with

0 < a < b < c < d < 8


1
Expert's answer
2021-12-07T06:49:15-0500

A Pythagorean quadruple is a tuple of integers "a, b, c," and "d," such that "a^2 + b^2 + c^2 = d^2."


"1^2+2^2+3^2=14, No"

"1^2+2^2+4^2=21, No"

"1^2+2^2+5^2=30, No"

"1^2+2^2+6^2=41, No"

"1^2+3^2+4^2=26, No"

"1^2+3^2+5^2=35, No"

"1^2+3^2+6^2=46, No"

"1^2+4^2+5^2=42, No"

"1^2+4^2+6^2=53, No"

"1^2+5^2+6^2=62, No"

"1^2+6^2+7^2=86, No"

"2^2+3^2+4^2=29, No"

"2^2+3^2+5^2=38, No"

"2^2+3^2+6^2=49=7^2, Yes"

"2^2+4^2+5^2=45, No"

"2^2+4^2+6^2=56, No"

"2^2+5^2+6^2=65, No"

"3^2+4^2+5^2=50, No"

"3^2+4^2+6^2=61, No"

"3^2+5^2+6^2=70, No"

"4^2+5^2+6^2=77, No"



"\{2,3,6,7\}"